# Predictive Control and Regret Analysis of Non-Stationary MDP with Look-ahead Information

### v1

@@ Appendix D Bounding the error of $Q$ function
$\leq$ $\epsilon_{0}+\delta_{0}D+2\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)+2\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right)$
where we used Proposition D.1. $\square$
We are now ready for Proof of Corollary 5.3
Proof:  [Proof of Corollary 5.3] Lemma D.3 shows that for any $s$, if the Proposed algorithm picks action $a$ instead of the optimal action $a^{*}$, we have
$(Q_{t}^{*}(s,a^{*})-\tilde{\Psi}_{t}(s,a^{*}))-(Q_{t}^{*}(s,a)-\tilde{\Psi}_{t}(s,a))\leq\gamma^{k/J}\operatorname{sp}\left(V_{t+k+1}^{*}\right)$
$\Rightarrow$ $(Q_{t}^{*}(s,a^{*})-\tilde{\Psi}_{t}(s,a^{*})+\hat{\Psi}_{t}(s,a^{*})-\hat{\Psi}_{t}(s,a^{*}))$
$-(Q_{t}^{*}(s,a)-\tilde{\Psi}_{t}(s,a)+\hat{\Psi}_{t}(s,a)-\hat{\Psi}_{t}(s,a))\leq\gamma^{k/J}\operatorname{sp}\left(V_{t+k+1}^{*}\right)$
$\Rightarrow$ $Q_{t}^{*}(s,a^{*})-Q_{t}^{*}(s,a)\leq\gamma^{k/J}\operatorname{sp}\left(V_{t+k+1}^{*}\right)+\underbrace{(\hat{\Psi}_{t}(s,a^{*})-\hat{\Psi}_{t}(s,a))}_{<0}$
$+\left|\tilde{\Psi}_{t}(s,a^{*})-\hat{\Psi}_{t}(s,a^{*})\right|+\left|\tilde{\Psi}_{t}(s,a)-\hat{\Psi}_{t}(s,a)\right|$
$\Rightarrow$ $Q_{t}^{*}(s,a^{*})-Q_{t}^{*}(s,a)\leq\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\left(V_{t+k+1}^{*}\right)+2\epsilon_{0}+2\delta_{0}D+4\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)$
@@
$\qquad\qquad+4\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right)$
where we used Lemma D.4 twice in the last inequality. $\square$
@@
We are now ready to prove Theorem 4.1.
@@
$\leq$ $\sum_{t=0}^{T-k-2}\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\left(V_{t+k+2}^{*}\right)+2T\epsilon_{0}+2T\delta_{0}D+4T\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)$ (29)
$+4T\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right)$ (30)
In (27) and (28), we expand out the value functions and rearrange the terms. Applying Corollary 5.3 leads to the resulting bound. $\square$
Appendix F Lower bound of regret for non-stationary MDP not satisfying 1
In order to show the necessity of 1, we provide the following counter-example that generates a linear regret for any $k$.
@@ Appendix F Lower bound of regret for non-stationary MDP not satisfying 1
For any fixed $k,T$ such that $k\ll T$, we can generate the following non-stationary MDP with three states $\{s^{i}\}_{i=1,2,3}$. The learner starts at $s^{1}$ and must make the decision of going to $s^{2}$ or $s^{3}$ at time step 00. Each of these actions has a reward of 00. $s^{2},s^{3}$ are sinks, and the learner can not move out after time step 1. Clearly, the above MDP does not satisfy 1, since for any $J>0$, $P_{t}^{\pi_{t}}\circ\cdots P_{t+J}^{\pi_{t+J}}(s^{1}|s^{2})=0$ for all policies $\boldsymbol{\pi}$. At time step 00, a state $s^{*}\in\{s_{2},s_{3}\}$ is chosen at random, such that the learner gets a reward of $1$ if at $s^{*}$ for all time steps after $k+1$, and a reward of 00 if otherwise. We claim that any algorithm with prediction horizon $k$ would generate a linear regret for the above MDP, as there is a non-zero constant probability for any policy to not choose $s^{*}$ at time step 00.

### v2

@@ Appendix D Bounding the error of $Q$ function
$\leq$ $2\epsilon_{0}+2\delta_{0}D+2\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)+2\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right),$
where we used Proposition D.1 in the last inequality. $\square$
We are now ready for Proof of Corollary˜5.3.
Proof: [Proof of Corollary˜5.3] Fix a state $s$ and time $t$. Let $a^{*}$ denote an optimal action at $(t,s)$, and let $a$ be the action chosen by Algorithm˜1. Define the error vector over actions
$e_{t}(a^{\prime}):=\hat{\Psi}_{t}(s,a^{\prime})-Q_{t}^{*}(s,a^{\prime}),\qquad a^{\prime}\in\mathcal{A}.$
We first relate the suboptimality gap to the span of this error:
$Q_{t}^{*}(s,a^{*})-Q_{t}^{*}(s,a)$ $=\hat{\Psi}_{t}(s,a^{*})-e_{t}(a^{*})-\bigl(\hat{\Psi}_{t}(s,a)-e_{t}(a)\bigr)$
$=\bigl(\hat{\Psi}_{t}(s,a^{*})-\hat{\Psi}_{t}(s,a)\bigr)+\bigl(e_{t}(a)-e_{t}(a^{*})\bigr).$
Since $a$ maximizes $\hat{\Psi}_{t}(s,\cdot)$, we have $\hat{\Psi}_{t}(s,a)\geq\hat{\Psi}_{t}(s,a^{*})$ and hence $\hat{\Psi}_{t}(s,a^{*})-\hat{\Psi}_{t}(s,a)\leq 0$. Therefore
$Q_{t}^{*}(s,a^{*})-Q_{t}^{*}(s,a)$ $\leq e_{t}(a)-e_{t}(a^{*})$
$\leq\max_{a^{\prime},a^{\prime\prime}}\bigl(e_{t}(a^{\prime})-e_{t}(a^{\prime\prime})\bigr)=\operatorname{sp}\bigl(e_{t}\bigr).$
That is,
$Q_{t}^{*}(s,a^{*})-Q_{t}^{*}(s,a)\;\leq\;\operatorname{sp}\bigl(\hat{\Psi}_{t}(s,\cdot)-Q_{t}^{*}(s,\cdot)\bigr).$ (27)
Next, use the triangle inequality for the span seminorm (Proposition˜A.1):
$\operatorname{sp}\bigl(\hat{\Psi}_{t}(s,\cdot)-Q_{t}^{*}(s,\cdot)\bigr)$ $\leq\operatorname{sp}\bigl(\hat{\Psi}_{t}(s,\cdot)-\tilde{\Psi}_{t}(s,\cdot)\bigr)+\operatorname{sp}\bigl(\tilde{\Psi}_{t}(s,\cdot)-Q_{t}^{*}(s,\cdot)\bigr).$ (28)
The second term in equation 28 is controlled by Lemma˜D.3:
$\operatorname{sp}\bigl(\tilde{\Psi}_{t}(s,\cdot)-Q_{t}^{*}(s,\cdot)\bigr)=\operatorname{sp}\bigl(Q_{t}^{*}(s,\cdot)-\tilde{\Psi}_{t}(s,\cdot)\bigr)\leq\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\bigl(V^{*}_{t+k+1}\bigr).$ (29)
The first term in equation 28 is bounded by Lemma D.4:
$\operatorname{sp}\bigl(\hat{\Psi}_{t}(s,\cdot)-\tilde{\Psi}_{t}(s,\cdot)\bigr)$ $\leq 2\epsilon_{0}+2\delta_{0}D$
$\quad+2\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\biggl(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}\,D\biggr)$
@@
$\quad+2\gamma^{\lfloor k/J\rfloor}\biggl(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor J+j}D\biggr).$ (30)
Combining equation 27, equation 28, equation 29 and equation 30, we obtain
$Q_{t}^{*}(s,a^{*})-Q_{t}^{*}(s,a)$ $\leq\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\bigl(V^{*}_{t+k+1}\bigr)+2\epsilon_{0}+2\delta_{0}D$
$\quad+2\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\biggl(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}\,D\biggr)$
$\quad+2\gamma^{\lfloor k/J\rfloor}\biggl(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor J+j}\,D\biggr).$
$\square$
@@
We are now ready to prove Theorem˜4.1.
@@
$\leq$ $\sum_{t=0}^{T-k-2}\gamma^{\lfloor k/J\rfloor}\operatorname{sp}\left(V_{t+k+2}^{*}\right)+2T\epsilon_{0}+2T\delta_{0}D+2T\sum_{i=0}^{\lceil k/J\rceil-1}\gamma^{i}\left(\sum_{j=1}^{J}\epsilon_{iJ+j}+\sum_{j=1}^{J}\delta_{iJ+j}D\right)$ (33)
$+2T\gamma^{\lfloor k/J\rfloor}\left(\sum_{j=1}^{k\%J}\epsilon_{\lfloor k/J\rfloor\cdot J+j}+\sum_{j=1}^{k\%J}\delta_{\lfloor k/J\rfloor\cdot J+j}D\right).$ (34)
In equation 31 and equation 32, we expand out the value functions and rearrange the terms. Applying Corollary˜5.3 leads to the resulting bound. $\square$
Appendix F Lower bound of regret for non-stationary MDP not satisfying ˜2
In order to show the necessity of ˜2, we provide the following counter-example that generates a linear regret for any $k$.
@@ Appendix F Lower bound of regret for non-stationary MDP not satisfying ˜2
For any fixed $k,T$ such that $k\ll T$, we can generate the following non-stationary MDP with three states $\{s^{i}\}_{i=1,2,3}$. The learner starts at $s^{1}$ and must make the decision of going to $s^{2}$ or $s^{3}$ at time step 0. Each of these actions has a reward of 0. $s^{2},s^{3}$ are sinks, and the learner can not move out after time step 1. Clearly, the above MDP does not satisfy ˜2, since for any $J>0$, $P_{t}^{\pi_{t}}\circ\cdots P_{t+J}^{\pi_{t+J}}(s^{1}|s^{2})=0$ for all policies $\bm{\pi}$. At time step 0, a state $s^{*}\in\{s_{2},s_{3}\}$ is chosen at random, such that the learner gets a reward of $1$ if at $s^{*}$ for all time steps after $k+1$, and a reward of 0 if otherwise. We claim that any algorithm with prediction horizon $k$ would generate a linear regret for the above MDP, as there is a non-zero constant probability for any policy to not choose $s^{*}$ at time step 0.